Find the measure of the indicated angles complementary angles with measures
A. 4x-9 B. 6x-1
Please describe how you solved the above problem. I am trying to learn how to solve this type of problem. I have been unsuccessful searching online. Thank you.

Found 2 solutions by oberobic, lwsshak3:
Answer by oberobic(2304)

Complementary angles total 90 degrees.
If one angle is 55, then the complementary angle = 90-55 = 35.
.
If you have one angle:
4x-9
then the complementary angle is:
90 - (4x-9)
= 81 -4x
.
But without additional information, it is impossible to tell you what the value of 'x' is.
.

Answer by lwsshak3(11628)

Find the measure of the indicated angles complementary angles with measures
A. 4x-9 B. 6x-1
**
let A= one of the complementary angles=4x-9
let B=the other complementary angle=6x-1
The sum of complementary angles=90º
A+B=90º
4x-9+6x-1=90
10x-10=90
10x=90+10=100
x=10º
complementary angle=90-10=80º
ans:
The two compementary angles are: 10º and 80º